For The Equation y=2(x)+15,and y=4x+7. Find an x-value that gives the same y-value for both equations
Answer by checkley77(12844)

y=2x+15
y=4x+7 SUBTRACT.
-----------------
0=-2X+8
2X=8
X=8/2
X=4 ANS.
PROOF:
Y=2*4+15
Y=8+15
Y=23
Y=4*4+7
Y=16+7
Y=23